# Primordial gravitational waves in Wheeler-DeWitt non-commutative linearized branch-cut quantum gravity

### 12 Jan 2024 v1

@@ Abstract
Branch-cut gravity (BCG) is an extended version of the ontological domain of General Relativity, which is analytically continued to the complex plane. When combined with the Hawking-Hertog multiverse conception, BCG successfully addresses the issue of the primordial singularity. It consistently portrays the early Universe as a Riemannian foliation in which the singularities of the multiverse merge, giving rise to a smooth branching topological structure that resembles continuously connected Riemann surfaces. This structure introduces a new cosmic scale factor that is analytically continued into the complex plane. In this contribution, we start with the recently developed Wheeler DeWitt-Hořava-Lifshitz non-commutative BCG formulation of quantum gravity. We investigate the impact of a non-commutative mini-superspace of variables obeying Poisson algebra on the accelerated behavior of the branch-cutting cosmic scale factor. Furthermore, we conduct a preliminary study on relic gravitational waves based on a linearized BCG approach, which describes the dynamics of a slightly perturbed gravitational field. Our findings concerning the dynamics of the scale factor reveal that they are highly sensitive to both the initial conditions and the different parameterizations of the primordial matter content. These findings indicate significant acceleration effects when compared to predictions based on conventional commutative approaches. These effects drive spacetime acceleration, offering a new perspective on explaining the accelerating expansion rate of our Universe. As far as primordial relic gravitational waves are concerned, our predictions reveal an intricate transition between the two phases of the branched Universe: a contraction phase preceding the conventional concept of a primordial singularity and a subsequent expansion phase whose transition region is characterized by a Riemannian topological foliation structure. Furthermore, this transition is characterized by asymmetric distributions of gravitational wave intensities. These results indicate in turn dynamic asymmetries in the distribution of matter and energy, arising from the capture of short and long scales within the non-commutative algebraic structure. These results are consistent with the expected highly inhomogeneous distribution of primordial gravitational waves, in contrast to the anticipated future observations of relic gravitational waves, which are expected to fill the universe as a stochastic homogeneous background.
preprint: APS/123-QED
I Introduction
@@ II Branch-cut Gravity
These findings expanded our understanding that enhancing the descriptive domain of a theory by incorporating complex or pseudo-complex variables can lead to a broader perspective on both infinitesimally small and large scales, with direct and tangible physical implications.
BGC extends GR to the complex domain, whence one obtains analytically continued Friedmann equations. These new Friedmann equations are then superposed, i.e., the equations are summed over some index $\xi$ which is associated to a set of independent universes, following the concept of multiverse proposed by Hawking-Hertog [29]. The Riemann sum is then taken to the integral limit. The resulting set of integro-differential equations - the BCG Friedmann equations - then leads to a new a scale parameter, $\ln^{-1}[\beta(t)]$, where $\beta(t)$ represents a regularization function. BGC thus provides a possible solution to the issue of a primordial singularity, by depicting the initial phases of the Universe as a set of Riemannian foliations. In other words, in the BCG formulation, the multiverse singularities merge, generating a smooth branching topological structure characterized by continuously connected manifolds.
@@
$ds_{[\rm{ac}]}^{2}=-\sigma^{2}N^{2}(t)c^{2}dt^{2}+\sigma^{2}\bigl{(}\ln^{-1}[\beta(t)]\bigr{)}^{2}\Biggl{[}\frac{dr^{2}}{\bigl{(}1-kr^{2}(t)\bigr{)}}+r^{2}(t)\Bigl{(}d\theta^{2}+\sin^{2}\theta d\phi^{2}\Bigr{)}\Biggr{]}\,.$ (1)
In this expression, $[\rm{ac}]$ denotes analytical continuation to the complex plane, the variables $r$ and $t$ represent real and complex space-time parameters, respectively, and $k$ denotes the spatial curvature of the multiverse, corresponding to negative curvature ($k=-1$), flat ($k=0$), or positively curved spatial hypersurfaces ($k=1$). The quantity $N(t)$ represents the lapse function, with $\sigma^{2}=2/3\pi$ denoting a normalization factor. The lapse function $N(t)$ is not dynamical but rather a pure gauge variable. Gauge invariance of the action in General Relativity yields a Hamiltonian constraint that requires a gauge-fixing condition on the lapse (see [30]).
The mapping between the scale parameter of the standard formulation, denoted as $a(t)$, and the corresponding scale factor in BCG, $\ln^{-1}(\beta(t))$, $a(t)\mapsto\ln^{-1}(\beta(t))$ does not correspond to a simple direct scale-factor parametrization based on the real nFriedman-Lemaître-Robertson-Walker (FLRW) single-pole metric, nor does it constitute a direct generalization of Friedmann’s equations [31]. Due to the non-linearity of Einstein’s field equations, such a direct generalization would formally not be consistent. The branch-cut formulation, on the other hand, arises from the complexification of the FLRW metric [31, 32, 33, 34], resulting in a sum of field equations associated to continuously distributed single-poles arranged along a line in the complex plane, each with infinitesimal residues (for details, see refs. [12, 13, 14, 15, 16, 17, 18, 8, 9]). Through a Riemann integration process, this complexification gives rise to the new scale factor, denoted as $\ln^{-1}(\beta(t))$, and a topological foliated spacetime structure. Branch-cut gravitation provides a gateway to the evolutionary phase preceding the primordial singularity, often referred to as a mirror world. In both its classical and quantum versions, this concept replaces the singularity with a topological transition between the contraction and expansion phases of the evolutionary universe. During this transition, spacetime takes on a helix-like topological shape, as proposed by branch-cut cosmology around a branch-point. Importantly, this process preserves the fundamental conservation laws of thermodynamics.
@@
III Non-commutative Branch-cut Quantum Gravity
The starting point of this study is a recently developed formulation based on the Wheeler-DeWitt-Hořava-Lifshitz quantum gravity [8, 9]. Unlike General Relativity, Hořava-Lifshitz theory is a renormalizable theory of gravity that retains Lorentz invariance at low energy while breaking this symmetry at high energy, which is a consequence of the implicit presence of a minimal length [11, 38]. The Hořava-Lifshitz approach incorporates high-order curvature contributions into the Lagrangian density, preserving the diffeomorphism property of General Relativity [39], an isomorphism of smooth manifolds, as well as the usual foliation of the Arnowitt-Deser-Misner (ADM) formalism at the Infrared Limit [40]. The Hořava-Lifshitz formulation, in combination with the Wheeler-DeWitt equation (WdW) [10], is ghost-free and thus suitable for describing quantum effects of the gravitational field [40]. The solutions of the WdW equation, represented by a geometric functional of compact manifolds and matter fields, describe the evolution of the quantum wave function of the Universe [41, 42]. An intriguing aspect of the WdW equation is the absence of a time variable, although it is linked to a classical image of space-time [39]. This feature is characteristic of the classical Hamilton-Jacobi formulation of General Relativity, where the observable Universe does not exhibit time-reversal symmetry, giving rise to a quantum description of events involving cosmic constituents [43, 44, 45, 46]. The dependence of the quantum wave function of the Universe on a “3-geometry”, corresponds to the equivalence class of metrics under a diffeomorphism, independent of the specific coordinates of the metric tensor [44, 45, 47]. As for the quantum cosmological interpretation of the wave function of the Universe, the prevailing view involves a functional restricted to a superspace configuration, encompassing three surfaces and matter fields denoted generically by $\Phi$. In this structure, the metric is represented by $h_{ij}$. The associated WdW wave function, denoted as $\Psi(h_{ij},\Phi)$, can be understood as a description of the evolution of $\Psi(\Phi)$ with respect to the physical variable $\Phi$.
@@ III Non-commutative Branch-cut Quantum Gravity
${\cal S}_{HL}$ $=$ $\frac{M_{P}^{2}}{2}\int d^{3}x\,dt\,N\sqrt{g}\Biggl{(}K_{ij}K^{ij}-\lambda K^{2}-g_{0}M^{2}_{P}-g_{1}{\cal R}-g_{2}M^{-2}_{P}{\cal R}^{2}-g_{3}M_{P}^{-2}{\cal R}_{ij}{\cal R}^{ij}--g_{4}M^{-4}_{P}{\cal R}^{3}$ (2)
-- $g_{5}M_{P}^{-4}{\cal R}({\cal R}^{i}_{j}{\cal R}^{j}_{i})-g_{6}M_{P}^{-4}{\cal R}^{i}_{j}{\cal R}^{j}_{k}{\cal R}^{k}_{i}-g_{7}M^{-4}_{P}{\cal R}\nabla^{2}{\cal R}\!-g_{8}M^{-4}_{P}\nabla_{i}{\cal R}_{jk}\nabla^{i}R^{jk}\Biggr{)}\,.$
In the BCG formulation, ${\cal S}_{HL}$ depends on the branching scalar curvature of the Universe, ${\cal R}$ and its derivatives, in different orders [11, 38, 40, 48, 18, 8, 9, 49, 50]. In expression (2), $g_{i}$ denote running coupling constants, $M_{P}$ is the Planck mass, $\nabla_{i}$ represents covariant derivatives, and the branching Ricci components of the three dimensional metrics may be determined by imposing a maximum symmetric surface foliation [18] which gives:
${\cal R}_{ij}=\frac{2}{\sigma^{2}u^{2}(t)}g_{ij}\,,\quad\mbox{and}\quad{\cal R}=\frac{6}{\sigma^{2}u^{2}(t)}\,,$ (3)
where the variable change $u(t)\equiv\ln^{-1}[\beta(t)]$, with $du\equiv d\ln^{-1}[\beta(t)]$, was introduced. $K=K^{ij}g_{ij}$ represents in expression (2) the trace of the extrinsic curvature tensor $K_{ij}$ [18, 8, 9]:
$K=K^{ij}g_{ij}=-\frac{3}{2\sigma Nu(t)}\frac{du(t)}{dt}.$ (4)
@@
${\cal H}=\frac{1}{2}\frac{N}{u}\Bigl{[}-p^{2}_{u}+g_{r}-g_{k}u^{2}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}\Bigr{]},$ (5)
@@
$g_{s}$$\equiv$ $288\pi^{4}\bigl{(}3\lambda-1\bigr{)}(9g_{4}+3g_{5}+g_{6}).$ (6)
@@
${\cal H}=\frac{1}{2}\frac{N}{u}\Bigl{[}-p^{2}_{u}+g_{r}-g_{m}u-g_{k}u^{2}-g_{q}u^{3}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}\Bigr{]}\,.$ (7)
@@
We then introduce a mapping which relates the commutative $\{\tilde{u},\tilde{p}_{u},\tilde{v},\tilde{p}_{v}\}$ and the non-commutative phase space set of variables $\{{u},{p}_{u},{v},{p}_{v}\}$ [9] (for the details, see Appendix: $\tilde{p}_{u}=(p_{u}-\chi p_{v});~{}\tilde{p}_{v}=(-\gamma p_{u}+\alpha u-\alpha v+p_{v})$). By inverting these equation, using (7) and the contribution of the perfect fluid, mentioned above (for details, please consult [9]), we arrive at the following expression for the super-Hamiltonian results [9]
${\cal H}=\frac{1}{2}\frac{N}{u}\Biggl{[}-\Bigl{(}\tilde{p}_{u}-\chi\tilde{p}_{v}\Bigr{)}^{2}-\frac{1}{u^{3\alpha-1}}\Bigl{(}\gamma\tilde{p}_{u}-\alpha u+\alpha v-\tilde{p}_{v}\Bigr{)}+\Biggl{(}g_{r}-g_{m}u-g_{k}u^{2}-g_{q}u^{3}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}\Biggr{)}\Biggr{]}\,.$ (9)
Canonical quantization procedures applied to the Hamiltonian (9), allow the variables $u(t)$ and $v(t)$ along with their corresponding conjugate momenta $p_{\rm{u}}$ and $p_{\rm{v}}$, to be treated as operators. Making the replacements $\tilde{p}_{u}\to-i\frac{\tilde{\partial}}{\partial u}$ and $\tilde{p}_{u}\to-i\frac{\tilde{\partial}}{\partial v}$, we obtain the following differential equation to describe the evolution of the wave function of the Universe in the non-commutative approach [9], giving the condition ${\cal H}\Psi(u,v)=0$:
$\frac{1}{2}\frac{N}{u}\Biggl{[}\Biggl{(}\frac{\tilde{\partial}^{2}}{\partial u^{2}}-2\chi\frac{\tilde{\partial}}{\partial u}\frac{\tilde{\partial}}{\partial v}+\chi^{2}\frac{\tilde{\partial}^{2}}{\partial v^{2}}\Biggr{)}+\frac{1}{u^{3\alpha-1}}\Biggl{(}i\gamma\frac{\tilde{\partial}}{\partial u}-i\frac{\tilde{\partial}}{\partial v}+\alpha u-\alpha v\Biggr{)}+\Biggl{(}g_{r}-g_{m}u-g_{k}u^{2}-g_{q}u^{3}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}\Biggr{)}\Biggr{]}\Psi(u,v)=0\,.$ (10)
@@
[(~$\Biggl{[}\Biggl{(}\frac{\tilde{\partial}^{2}}{\partial u^{2}}-2\chi\frac{\tilde{\partial}}{\partial u}\frac{\tilde{\partial}}{\partial v}+\chi^{2}\frac{\tilde{\partial}^{2}}{\partial v^{2}}\Biggr{)}-\frac{1}{u^{3\alpha-1}}\Biggl{(}|\gamma|\frac{\tilde{\partial}}{\partial u}+i\frac{\tilde{\partial}}{\partial v}-\alpha u+\alpha v\Biggr{)}+\Biggl{(}g_{r}-g_{m}u-g_{k}u^{2}-g_{q}u^{3}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}\Biggr{)}\Biggr{]}\Psi(u,v)=0\,.$ (11)
For simplicity, in the following we do not use the symbol tilde in the partial derivatives as identification of the new variables in the scope of non-commutative algebra. Equation (11) can be rewritten in the general form:
$\Biggl{[}a(u,v)\frac{\partial^{2}}{\partial u^{2}}-2b(u,v)\frac{\partial^{2}}{\partial u\partial v}+c(u,v)\frac{\partial^{2}}{\partial v^{2}}\Biggr{]}\Psi(u,v)=\Biggl{[}d(u,v)\frac{\partial}{\partial u}+e(u,v)\frac{\partial}{\partial v}+F\bigl{(}u,v\bigr{)}\Biggr{]}\Psi(u,v)\,,=G\Biggl{(}u,v,\Psi,\frac{\partial}{\partial u},\frac{\partial}{\partial v}\Biggr{)}\,,$ (12)
@@
$F\bigl{(}u,v\bigr{)}$ $=$ $-\Bigl{(}g_{r}-g_{m}u-g_{k}u^{2}-g_{q}u^{3}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}-\frac{\alpha}{u^{3\alpha-2}}+\frac{\alpha v}{u^{3\alpha-1}}\Bigr{)}\,,$ (13)
where $a(u,v)$, $b(u,v)$, and $c(u,v)$ represent functions of the independent variables $u$ and $v$, and have continuous derivatives up to the second-order. Since the $b^{2}(u,v)-a(u,v)c(u,v)=0$ expression (12) belongs to the mathematical group of parabolic equations, to reduce this equation to a canonical form, one should first write out the characteristic equation [58]
$a~{}du^{2}-2b~{}du~{}dv+c~{}dv^{2}=0\,,$ (14)
@@
$a~{}d\eta-\bigl{(}b\pm\sqrt{b^{2}-ac}\bigr{)}d\xi=0\,.$ (15)
Then, one should find their general integrals. In the case of a parabolic equation, the two previous solutions coincide, resulting in a common general integral $\varphi(u,v)={\cal I}_{G}.$ This allows the variables $u$ and $v$ to be changed to new independent variables $\xi$ and $\eta$, in accordance with
$\xi=\varphi(u,v),\quad\mbox{and}\quad\eta=\eta(u,v)\,,$ (16)
where $\eta(u,v)$ is a differentiable function that satisfies the non-degeneracy condition of the Jacobian $D(\xi,\eta)/D(u,v)$ in the given domain. As a result, Equation (12) is reduced to the canonical form
$\frac{\partial^{2}\Psi(\xi,\eta)}{\partial\eta^{2}}=G\Biggl{(}\xi,\eta,\Psi,\frac{\partial}{\partial\xi},\frac{\partial}{\partial\eta}\Biggr{)}\,.$ (17)

### 03 Feb 2024 v2

# Dynamical equations and primordial GWs in a non-commutative branching cosmology

@@ Abstract
Branch-cut Gravity (BCG) an extended version of the ontological cosmology domain of General Relativity, which is analytically continued to the complex plane. When combined with the Hawking-Hertog multiverse conception, BCG successfully addresses the issue of the primordial singularity. It consistently portrays the early Universe as a Riemannian foliation in which the singularities of the multiverse merge, giving rise to a smooth branching topological structure that resembles continuously connected Riemann surfaces. This structure introduces a new cosmic scale factor that is analytically continued into the complex plane. In this contribution, we start with the recently developed Wheeler DeWitt-Hořava-Lifshitz non-commutative BCG formulation of quantum gravity. We investigate the impact of a non-commutative mini-superspace of variables obeying Poisson algebra on the accelerated behavior of the branch-cutting cosmic scale factor. Furthermore, we conduct a preliminary study on relic gravitational waves based on a linearized BCG approach, which describes the dynamics of a slightly perturbed gravitational field. Our findings concerning the dynamics of the scale factor reveal that they are highly sensitive to both the initial conditions and the different parameterizations of the primordial matter content. These findings indicate significant acceleration effects when compared to predictions based on conventional commutative approaches. These effects drive spacetime acceleration, offering a new perspective on explaining the accelerating expansion rate of our Universe. As far as primordial relic gravitational waves are concerned, our predictions reveal an intricate transition between the two phases of the branched Universe: a contraction phase preceding the conventional concept of a primordial singularity and a subsequent expansion phase whose transition region is characterized by a Riemannian topological foliation structure. Furthermore, this transition is characterized by asymmetric distributions of gravitational wave intensities. These results indicate in turn dynamic asymmetries in the distribution of matter and energy, arising from the capture of short and long scales within the non-commutative algebraic structure. These results are consistent with the expected highly inhomogeneous distribution of primordial gravitational waves, in contrast to the anticipated future observations of relic gravitational waves, which are expected to fill the universe as a stochastic homogeneous background.
1 Introduction
@@ 2 Branch-cut Gravity
These findings expanded our understanding that enhancing the descriptive domain of a theory by incorporating complex or pseudo-complex variables can lead to a broader perspective on both infinitesimally small and large scales, with direct and tangible physical implications. BGC extends GR to the complex domain, whence one obtains analytically continued Friedmann equations. These new Friedmann equations are then superposed, i.e., the equations are summed over some index $\xi$ which is associated to a set of independent universes, following the concept of multiverse proposed by Hawking-Hertog [29]. The Riemann sum is then taken to the integral limit. The resulting set of integro-differential equations - the BCG Friedmann equations - then leads to a new a scale parameter, $\ln^{-1}[\beta(t)]$, where $\beta(t)$ represents a regularization function. BGC thus provides a possible solution to the issue of a primordial singularity, by depicting the initial phases of the Universe as a set of Riemannian foliations. In other words, in the BCG formulation, the multiverse singularities merge, generating a smooth branching topological structure characterized by continuously connected manifolds.
@@
$ds_{[\rm{ac}]}^{2}=-\sigma^{2}N^{2}(t)c^{2}dt^{2}+\sigma^{2}\bigl{(}\ln^{-1}[\beta(t)]\bigr{)}^{2}\Biggl{[}\frac{dr^{2}}{\bigl{(}1-kr^{2}(t)\bigr{)}}+r^{2}(t)\Bigl{(}d\theta^{2}+\sin^{2}\theta d\phi^{2}\Bigr{)}\Biggr{]}\,.$ (2.1)
In this expression, $[\rm{ac}]$ denotes analytical continuation to the complex plane, the variables $r$ and $t$ represent real and complex space-time parameters, respectively, and $k$ denotes the spatial curvature of the multiverse, corresponding to negative curvature ($k=-1$), flat ($k=0$), or positively curved spatial hypersurfaces ($k=1$). The quantity $N(t)$ represents the lapse function, with $\sigma^{2}=2/3\pi$ denoting a normalization factor. The lapse function $N(t)$ is not dynamical but rather a pure gauge variable. Gauge invariance of the action in General Relativity yields a Hamiltonian constraint that requires a gauge-fixing condition on the lapse (see [30]). The mapping between the scale parameter of the standard formulation, denoted as $a(t)$, and the corresponding scale factor in BCG, $\ln^{-1}(\beta(t))$, $a(t)\mapsto\ln^{-1}(\beta(t))$ does not correspond to a simple direct scale-factor parametrization based on the real nFriedman-Lemaître-Robertson-Walker (FLRW) single-pole metric, nor does it constitute a direct generalization of Friedmann’s equations [31]. Due to the non-linearity of Einstein’s field equations, such a direct generalization would formally not be consistent. The branch-cut formulation, on the other hand, arises from the complexification of the FLRW metric [31, 32, 33, 34], resulting in a sum of field equations associated to continuously distributed single-poles arranged along a line in the complex plane, each with infinitesimal residues (for details, see refs. [12, 13, 14, 15, 16, 17, 18, 8, 9]). Through a Riemann integration process, this complexification gives rise to the new scale factor, denoted as $\ln^{-1}(\beta(t))$, and a topological foliated spacetime structure. Branch-cut gravitation provides a gateway to the evolutionary phase preceding the primordial singularity, often referred to as a mirror world. In both its classical and quantum versions, this concept replaces the singularity with a topological transition between the contraction and expansion phases of the evolutionary universe. During this transition, spacetime takes on a helix-like topological shape, as proposed by branch-cut cosmology around a branch-point. Importantly, this process preserves the fundamental conservation laws of thermodynamics.
@@
3 Non-commutative Branch-cut Quantum Gravity
The starting point of this study is a recently developed formulation based on the Wheeler-DeWitt-Hořava-Lifshitz quantum gravity [8, 9]. Unlike General Relativity, Hořava-Lifshitz theory is a renormalizable theory of gravity that retains Lorentz invariance at low energy while breaking this symmetry at high energy, which is a consequence of the implicit presence of a minimal length [11, 38]. The Hořava-Lifshitz approach incorporates high-order curvature contributions into the Lagrangian density, preserving the diffeomorphism property of General Relativity [39], an isomorphism of smooth manifolds, as well as the usual foliation of the Arnowitt-Deser-Misner (ADM) formalism at the Infrared Limit [40]. The Hořava-Lifshitz formulation, in combination with the Wheeler-DeWitt equation (WdW) [10], is ghost-free and thus suitable for describing quantum effects of the gravitational field [40]. The solutions of the WdW equation, represented by a geometric functional of compact manifolds and matter fields, describe the evolution of the quantum wave function of the Universe [41, 42]. An intriguing aspect of the WdW equation is the absence of a time variable, although it is linked to a classical image of space-time [39]. This feature is characteristic of the classical Hamilton-Jacobi formulation of General Relativity, where the observable Universe does not exhibit time-reversal symmetry, giving rise to a quantum description of events involving cosmic constituents [43, 44, 45]. The dependence of the quantum wave function of the Universe on a “3-geometry”, corresponds to the equivalence class of metrics under a diffeomorphism, independent of the specific coordinates of the metric tensor [44, 45, 47]. As for the quantum cosmological interpretation of the wave function of the Universe, the prevailing view involves a functional restricted to a superspace configuration, encompassing three surfaces and matter fields denoted generically by $\Phi$. In this structure, the metric is represented by $h_{ij}$. The associated WdW wave function, denoted as $\Psi(h_{ij},\Phi)$, can be understood as a description of the evolution of $\Psi(\Phi)$ with respect to the physical variable $\Phi$.
@@ 3 Non-commutative Branch-cut Quantum Gravity
${\cal S}_{HL}$ $=$ $\frac{M_{P}^{2}}{2}\int d^{3}x\,dt\,N\sqrt{g}\Biggl{(}K_{ij}K^{ij}-\lambda K^{2}-g_{0}M^{2}_{P}-g_{1}{\cal R}$ (3.1)
$-g_{2}M^{-2}_{P}{\cal R}^{2}-g_{3}M_{P}^{-2}{\cal R}_{ij}{\cal R}^{ij}-g_{4}M^{-4}_{P}{\cal R}^{3}-g_{5}M_{P}^{-4}{\cal R}({\cal R}^{i}_{j}{\cal R}^{j}_{i})$
$-g_{6}M_{P}^{-4}{\cal R}^{i}_{j}{\cal R}^{j}_{k}{\cal R}^{k}_{i}-g_{7}M^{-4}_{P}{\cal R}\nabla^{2}{\cal R}\!-g_{8}M^{-4}_{P}\nabla_{i}{\cal R}_{jk}\nabla^{i}R^{jk}\Biggr{)}\,.$
In the BCG formulation, ${\cal S}_{HL}$ depends on the branching scalar curvature of the Universe, ${\cal R}$ and on its derivatives, in different orders [11, 38, 40, 48, 18, 8, 9, 49, 50]. In expression (3.1), $g_{i}$ denote running coupling constants, $M_{P}$ is the Planck mass, $\nabla_{i}$ represents covariant derivatives, and the branching Ricci components of the three dimensional metrics may be determined by imposing a maximum symmetric surface foliation [18] which gives:
${\cal R}_{ij}=\frac{2}{\sigma^{2}u^{2}(t)}g_{ij}\,,\quad\mbox{and}\quad{\cal R}=\frac{6}{\sigma^{2}u^{2}(t)}\,,$ (3.2)
where the variable change $u(t)\equiv\ln^{-1}[\beta(t)]$, with $du\equiv d\ln^{-1}[\beta(t)]$, was introduced. $K=K^{ij}g_{ij}$ represents in expression (3.1) the trace of the extrinsic curvature tensor $K_{ij}$ [18, 8, 9]:
$K=K^{ij}g_{ij}=-\frac{3}{2\sigma Nu(t)}\frac{du(t)}{dt}.$ (3.3)
@@
${\cal H}=\frac{1}{2}\frac{N}{u}\Bigl{[}-p^{2}_{u}+g_{r}-g_{k}u^{2}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}\Bigr{]},$ (3.4)
@@
$g_{s}$$\equiv$ $288\pi^{4}\bigl{(}3\lambda-1\bigr{)}(9g_{4}+3g_{5}+g_{6}).$ (3.5)
@@
${\cal H}=\frac{1}{2}\frac{N}{u}\Bigl{[}-p^{2}_{u}+g_{r}-g_{m}u-g_{k}u^{2}-g_{q}u^{3}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}\Bigr{]}\,.$ (3.6)
@@
We then introduce a mapping which relates the commutative $\{\tilde{u},\tilde{p}_{u},\tilde{v},\tilde{p}_{v}\}$ and the non-commutative phase space set of variables $\{{u},{p}_{u},{v},{p}_{v}\}$ [9] (for the details, see Appendix: $\tilde{p}_{u}=(p_{u}-\chi p_{v});~{}\tilde{p}_{v}=(-\gamma p_{u}+\alpha u-\alpha v+p_{v})$). By inverting these equation, using (3.6) and the contribution of the perfect fluid, mentioned above (for details, please consult [9]), we arrive at the following expression for the super-Hamiltonian results [9]
${\cal H}$ $=$ $\frac{1}{2}\frac{N}{u}\Biggl{[}-\Bigl{(}\tilde{p}_{u}-\chi\tilde{p}_{v}\Bigr{)}^{2}-\frac{1}{u^{3\alpha-1}}\Bigl{(}\gamma\tilde{p}_{u}-\alpha u+\alpha v-\tilde{p}_{v}\Bigr{)}$ (3.8)
$+\Biggl{(}g_{r}-g_{m}u-g_{k}u^{2}-g_{q}u^{3}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}\Biggr{)}\Biggr{]}\,.$
Canonical quantization procedures applied to the Hamiltonian (3.8), allow the variables $u(t)$ and $v(t)$ along with their corresponding conjugate momenta $p_{\rm{u}}$ and $p_{\rm{v}}$, to be treated as operators. Making the replacements $\tilde{p}_{u}\to-i\frac{\tilde{\partial}}{\partial u}$ and $\tilde{p}_{u}\to-i\frac{\tilde{\partial}}{\partial v}$, we obtain the following differential equation to describe the evolution of the wave function of the Universe in the non-commutative approach [9], giving the condition ${\cal H}\Psi(u,v)=0$:
${\cal H}\Psi(u,v)$ $=$ $\frac{1}{2}\frac{N}{u}\Biggl{[}\Biggl{(}\frac{\tilde{\partial}^{2}}{\partial u^{2}}-2\chi\frac{\tilde{\partial}}{\partial u}\frac{\tilde{\partial}}{\partial v}+\chi^{2}\frac{\tilde{\partial}^{2}}{\partial v^{2}}\Biggr{)}+\frac{1}{u^{3\alpha-1}}\Biggl{(}i\gamma\frac{\tilde{\partial}}{\partial u}-i\frac{\tilde{\partial}}{\partial v}+\alpha u-\alpha v\Biggr{)}$ (3.9)
$+\Biggl{(}g_{r}-g_{m}u-g_{k}u^{2}-g_{q}u^{3}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}\Biggr{)}\Biggr{]}\Psi(u,v)=0\,.$
@@
${\cal H}\Psi(u,v)$ $=$ [(~$\Biggl{[}\Biggl{(}\frac{\tilde{\partial}^{2}}{\partial u^{2}}-2\chi\frac{\tilde{\partial}}{\partial u}\frac{\tilde{\partial}}{\partial v}+\chi^{2}\frac{\tilde{\partial}^{2}}{\partial v^{2}}\Biggr{)}-\frac{1}{u^{3\alpha-1}}\Biggl{(}|\gamma|\frac{\tilde{\partial}}{\partial u}+i\frac{\tilde{\partial}}{\partial v}-\alpha u+\alpha v\Biggr{)}$ (3.10)
$+\Biggl{(}g_{r}-g_{m}u-g_{k}u^{2}-g_{q}u^{3}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}\Biggr{)}\Biggr{]}\Psi(u,v)=0\,.$
For simplicity, in the following we do not use the symbol tilde in the partial derivatives as identification of the new variables in the scope of non-commutative algebra. Equation (3.10) can be rewritten in the general form:
$\Biggl{\{}\Biggl{[}\!a(u,v)\frac{\partial^{2}}{\partial u^{2}}-2b(u,v)\frac{\partial^{2}}{\partial u\partial v}+c(u,v)\frac{\partial^{2}}{\partial v^{2}}\!\Biggr{]}-\Biggl{[}\!d(u,v)\frac{\partial}{\partial u}+e(u,v)\frac{\partial}{\partial v}+F\bigl{(}u,v\bigr{)}\!\Biggr{]}\!\Biggr{\}}\Psi(u,v)=0$ (3.11)
@@
$F\bigl{(}u,v\bigr{)}$ $=$ $-\Bigl{(}g_{r}-g_{m}u-g_{k}u^{2}-g_{q}u^{3}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}-\frac{\alpha}{u^{3\alpha-2}}+\frac{\alpha v}{u^{3\alpha-1}}\Bigr{)}\,,$ (3.12)
where $a(u,v)$, $b(u,v)$, and $c(u,v)$ represent functions of the independent variables $u$ and $v$, and have continuous derivatives up to the second-order. Since the $b^{2}(u,v)-a(u,v)c(u,v)=0$ expression (3.11) belongs to the mathematical group of parabolic equations. In order to reduce this equation to a canonical form, one should first write out the characteristic equation [58]
$a~{}du^{2}-2b~{}du~{}dv+c~{}dv^{2}=0\,,$ (3.13)
@@
$a~{}d\eta-\bigl{(}b\pm\sqrt{b^{2}-ac}\bigr{)}d\xi=0\,.$ (3.14)
And then, one should find their general integrals. In the case of a parabolic equation, the two previous solutions coincide, resulting in a common general integral $\varphi(u,v)={\cal I}_{G}.$ This allows the variables $u$ and $v$ to be changed to new independent variables $\xi$ and $\eta$, in accordance with
$\xi=\varphi(u,v),\quad\mbox{and}\quad\eta=\eta(u,v)\,,$ (3.15)
where $\eta(u,v)$ is a differentiable function that satisfies the non-degeneracy condition of the Jacobian $D(\xi,\eta)/D(u,v)$ in the given domain. As a result, Equation (3.11) is reduced to the canonical form
$\frac{\partial^{2}\Psi(\xi,\eta)}{\partial\eta^{2}}=G\Biggl{(}\xi,\eta,\Psi,\frac{\partial}{\partial\xi},\frac{\partial}{\partial\eta}\Biggr{)}\,.$ (3.16)